solve the following initial value problem using laplace transforms 2d^2y/dt^3+3dy/dt-2y=e^-2t y=0,y*=-2


1
Expert's answer
2021-02-24T06:08:15-0500

First apply the Laplace transform to both sides of the equation, we will note the Laplace transform of f(t) as F(s):

"2s^2F(s)-2sf(0)-2f'(0)+3sF(s)-3f(0)-2F(s)=\frac{1}{s+2}"

Now using the initial conditions we get:

"(2s^2+3s-2)F(s)+4=\frac{1}{s+2}"

"F(s) = \frac{-4}{2s^2+3s-2}+\frac{1}{(s+2)(2s^2+3s-2)}"

Now to find the inverse transform, we will use the linearity and we will decompose the fractions into the simple elements:

"2s^2+3s-2=(2s-1)(s+2)"

"\frac{-4}{2s^2+3s-2}=\frac{A}{2s-1}+\frac{B}{s+2}"

"\begin{cases} A+2B=0 \\ 2A-B=-4 \end{cases}" gives us "\begin{cases} A=-8/5 \\ B=4/5 \end{cases}"

"(s+2)(2s^2+3s-2)=(s+2)^2(2s-1)"

"\frac{1}{(s+2)^2(2s-1)}=\frac{Cs+D}{(s+2)^2}+\frac{E}{2s-1}"

"\begin{cases} 2C+E=0 \\ -C+4E+2D=0 \\ -D+4E=1 \end{cases}" gives us "\begin{cases} C = -2/25 \\ D=-9/25 \\ E= 4/25\end{cases}" , which we can write as

"\frac{-2}{25}\frac{1}{s+2}-\frac{1}{5}\frac{1}{(s+2)^2}+\frac{2}{25}\frac{1}{s-1/2}"

Now by applying the inverse transformation term by term we find:

"f(t)=-\frac{4}{5}e^{t/2}+\frac{4}{5}e^{-2t}-\frac{2}{25}e^{-2t}-\frac{1}{5}e^{-2t}t+\frac{2}{25}e^{t/2}"

"f(t)=-\frac{18}{25}e^{t/2}+(\frac{18}{25}-\frac{1}{5}t)e^{-2t}"